OpenStudy (anonymous):

there are three ants in three corners of a triangle? each ant started moving towards another corner what is probability that they wont collide at any place?? answer: they need to travel either clockwise or anticlockwise the probability is 1/4. my question is how this 1/4 came ? can any one explain??

OpenStudy (anonymous):

you have i think 8 possibilities for ants moving left and right yes?

OpenStudy (anonymous):

3 ants, to directions each and 2^3=8

OpenStudy (anonymous):

like saying "flip a coin, what is the probability you get all head or all tails?'

OpenStudy (anonymous):

sample space looks like ( R R R) (R R L) (R L R) (L R R) (L L R) (L R L) (R L L ) (R R R) with r and l being obviously left and right.

OpenStudy (anonymous):

only two of these has all ants going in same direction, namely ( R R R) and (L L L ) therefore 2/8=1/4 is the answer

OpenStudy (anonymous):

hope this makes some sense

OpenStudy (anonymous):

hey cool.very very thanks for brief and clear explanation. get an medal.. once again thanks....